# Bilinear Subspace Variational Bayesian Inference for Joint Scattering Environment Sensing and Data Recovery in ISAC Systems

### v1

@@ Abstract
ISAC, joint sensing and data recovery, blinear subspace variational Bayesian inference. Then the joint problem is formulated as a bilinear structured sparse
@@ I Introduction
Joint scattering environment sensing and channel estimation: Since multiple users in the same cell share a common scattering environment, the three-dimensional (3D) scatterers seen by different users have a partially overlapping structure. The work in [5, 6] revealed that MU channels exhibit a hidden joint sparsity in the angular-domain due to the shared scattering environment, which can be exploited to enhance the channel estimation performance. [3, 7] concentrated on the partially overlapping structure between radar targets and communication scatterers, and a joint target sensing and channel estimation scheme was designed based the joint sparsity of radar and communication channels. The above works only consider the oversimplified case when MU or ISAC channels share a partially overlapping support in the angular domain, while the shared location-domain information is not fully used. A recent work in [8, 9] has proposed a two-dimensional (2D) location-domain sparse channel model to exploit the partially overlapping support of MU channels for the special case when all scatterers lies in a 2D area. However, the 3D location-domain sparse channel modeling is still not fully investigated. Moreover, the existing works perform joint scattering environment sensing and channel estimation based on pilot signals only. Nevertheless, data signals also provides much information about wireless channels that can help reduce pilot overhead.
@@
  • BiSVBI algorithm: We model the joint sensing and data recovery problem as a bilinear structured sparse signal recovery problem with a dynamic grid and imperfect factors. Then, we propose an expectation maximization based turbo BiSVBI (EM-Turbo-BiSVBI) algorithm to solve this problem efficiently, where the E-step performs Bayesian estimation of the location-domain sparse channels and the M-step refines the dynamic grid and learns the imperfect factors via gradient update. In particular, the E-step is based on a novel BiSVBI algorithm by minimizing the KL-divergence (KLD) between the variational posterior distribution and the true posterior associated with the bilinear model. To reduce the complexity, an independent constraint is imposed on the variational posterior distribution of the sparse channels to avoid matrix inverse in the calculation of the posterior covariance matrix. Moreover, a subspace constrained matrix inverse followed by a gradient update is proposed to replace the high dimensional matrix inverse in the calculation of the posterior mean. Theoretical analysis and simulations show that BiSVBI greatly reduces the algorithm complexity with almost no sacrifice on the performance and convergence speed.
@@ II-B Uplink Signal Model for Joint Sensing and Data Recovery
where $\alpha_{k,0}$ is the complex gain of the LoS channel path, $\alpha_{k,l}$ is the complex gain of the $l$-th NLoS channel path, $\tau_{o,k}$ denotes the timing synchronization error between user $k$ and the BS, $\tau_{0}\left(\boldsymbol{p}_{u,k}\right)=\frac{r_{l}}{c}$ is the propagation delay of the LoS path from user $k$ to the BS with $c$ denoting the speed of light,
@@ III-B Three-layer BGG Prior Model
where $\delta\left(\cdot\right)$ is the Dirac Delta function, $\Gamma\left(\rho;a,b\right)$ is a Gamma hyper-prior with shape parameter $a$ and rate parameter $b$. When $s_{k,q}=1$, $\alpha_{k,q}$ is a non-zero element and the corresponding variance $1/\rho_{k,q}$ is $\Theta\left(1\right)$. In this case, $a$ and $b$ should be chosen to satisfy $\tfrac{a}{b}=\mathbb{E}\left(\rho_{k,q}\right)=\Theta\left(1\right)$. When $s_{k,q}=0$, $\alpha_{k,q}$ is a zero element and the corresponding variance $1/\rho_{k,q}$ is close to zero. In this case, $\overline{a}$ and $\overline{b}$ should be chosen to satisfy $\tfrac{\overline{a}}{\overline{b}}=\mathbb{E}\left(\rho_{k,q}\right)\gg 1$. A typical value is $a=1$, $b=1$, $\overline{a}=1$, and $\overline{b}=10^{-5}$ [20].
@@ III-C Bilinear Sparse Bayesian Inference with Uncertain Parameters
where $\overline{\boldsymbol{v}}\triangleq\left\{\boldsymbol{\alpha},\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{s},\overline{\boldsymbol{s}},\gamma\right\}$ denotes the collection of all variables, and $p\left(\boldsymbol{\xi}\right)$ denotes the known prior distribution of $\boldsymbol{\xi}$. Once we obtain the MAP estimate of $\boldsymbol{\xi}^{\ast}$, we can obtain the minimum mean square error (MMSE) estimate of $\boldsymbol{\alpha}$ as $\boldsymbol{\alpha}^{*}=\int_{\boldsymbol{\alpha}}\boldsymbol{\alpha}p\left(\boldsymbol{\alpha}\mid\boldsymbol{y};\boldsymbol{\xi}^{\ast}\right)$, and the MAP estimate of $\boldsymbol{s}$ as $s_{k,q}^{\ast}=\arg\max_{s_{k,q}}p\left(s_{k,q}\mid\boldsymbol{y};\boldsymbol{\xi}^{\ast}\right),\forall k,q$. Moreover, the marginal posteriors $p\left(\boldsymbol{x}\mid\boldsymbol{y};\boldsymbol{\xi}\right)$ for the data can be used as a soft information for digital demodulation and channel decoding.
@@ IV-B Turbo-BiSVBI-E Step
For each turbo iteration, Module A treats $\textrm{h}_{k,q}^{\text{A}}$ as the prior and performs the BiSVBI to calculate the approximate conditional posteriors $q\left(s_{k,q}\right)$’s. Then, Module A passes extrinsic messages to Module B by subtracting the prior information from posterior information, i.e.,
@@ IV-C1 Formulation of BiSVBI
The BiSVBI module generalizes the SC-VBI in [17] from a linear observation model to a bilinear model to calculate the approximate marginal posteriors $q\left(\boldsymbol{v}\right)$ for fixed sensing parameters $\boldsymbol{\xi}$, where $\boldsymbol{v}\triangleq\left\{\boldsymbol{\alpha},\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{s},\gamma\right\}$. To be more specific, the turbo-iteration factor nodes $\textrm{h}_{k,q}^{\text{A}}\left(s_{k,q}\right),\forall k,q$ incorporate the prior information from Module B. Therefore, the following prior distribution is assumed in the BiSVBI module:
@@ IV-C2 Derivation of the BiSVBI Algorithm for $\mathcal{\mathscr{A}}_{\mathrm{VBI}}$
where $\mathbf{W}_{\alpha}=\textrm{Diag}\left(\left\langle\boldsymbol{\rho}\right\rangle\right)+\left\langle\kappa\right\rangle\left\langle\mathbf{A}_{\alpha}^{H}\mathbf{A}_{\alpha}\right\rangle$, $\mathrm{W}_{\alpha,k,q}$ is the $\left[\left(k-1\right)\left(Q+1\right)+q+1\right]$-th diagonal element of $\mathbf{W}_{\alpha}$, and $\boldsymbol{b}=\left\langle\kappa\right\rangle\left\langle\mathbf{A}_{\alpha}^{H}\right\rangle\boldsymbol{y}$ with $\mathbf{A}_{\alpha}=\left[\mathbf{A}_{\alpha}^{p};\mathbf{A}_{\alpha}^{d}\right]$,
@@
where $\boldsymbol{\mu}_{x,t_{d},n}$ and $\boldsymbol{\sigma}_{x,t_{d},n}^{2}$ are the posterior mean and variance corresponding to $\boldsymbol{x}_{t_{d},n}$, $\mathbf{W}_{x,t_{d},n}\triangleq\textrm{Diag}\left(\boldsymbol{\sigma}_{t_{d},n}^{2}\right)+\left\langle\kappa\right\rangle\left\langle\mathbf{\Psi}_{n}^{H}\mathbf{\Psi}_{n}\right\rangle$ with $\boldsymbol{\sigma}_{t_{d},n}^{2}=\left[\sigma_{k,n}^{2}\right]_{k\in\mathcal{K}_{t_{d},n}^{d}}$, $\boldsymbol{b}_{x,t_{d},n}=\left\langle\kappa\right\rangle\left\langle\mathbf{\Psi}_{n}^{H}\right\rangle\boldsymbol{y}_{t_{d},n}^{d}$, and $\left[\mathrm{W}_{x,t_{d},k,n}\right]_{k\in\mathcal{K}_{t_{d},n}^{d}}=\textrm{Diag}\left(\mathbf{W}_{x,t_{d},n}\right)$.
Please refer to Appendix -A for the detailed proof. Based on (25) and Lemma 2, the update equations of all variables are given below.
@@
$\boldsymbol{\mu}_{\alpha,\hat{\mathcal{S}}}^{s}=\left\langle\kappa\right\rangle\mathbf{W}_{\alpha,\hat{\mathcal{S}}}^{-1}\left\langle\mathbf{A}_{\alpha,\hat{\mathcal{S}}}^{H}\right\rangle\boldsymbol{y},$ (29)
@@
To improve the robustness of the algorithm against the support estimation error, we choose the best initial point $\boldsymbol{\mu}_{\alpha}^{0}$ from the following two choices: 1) $\boldsymbol{\mu}_{\alpha}^{s}$ obtained by the subspace constrained matrix inverse as in (29); 2) the posterior mean $\boldsymbol{\mu}_{\alpha}^{p}$ from the previous iteration, depending on which choice gives a lower value of the objective function $\varphi\left(\boldsymbol{\mu}_{\alpha}^{0}\right)$. In the first few iterations, $\boldsymbol{\mu}_{\alpha}^{s}$ is expected to be a good choice because the subspace constrained matrix inverse can be used to accelerate the initial convergence speed, compared to the MM-based methods in [23, 9, 24] which completely avoid any matrix inverse. As the iterations go on, the posterior mean $\boldsymbol{\mu}_{\alpha}^{p}$ from the previous iteration becomes more and more accurate and it may also become a good initial point.
@@
$\boldsymbol{\mu}_{\alpha}^{\left(i\right)}$ $=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}-\epsilon_{\alpha}^{\left(i\right)}\nabla_{\boldsymbol{u}}\varphi\left(\boldsymbol{u}\right)\mid_{\boldsymbol{u}=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}},$ (30)
@@
$\nabla_{\boldsymbol{u}}\varphi\left(\boldsymbol{u}=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}\right)=\mathbf{W}_{\alpha}\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}-\boldsymbol{b}_{\alpha},$ (31)
@@
Once we obtain the estimated posterior mean $\boldsymbol{\mu}_{\alpha}$, we can obtain an estimated support $\hat{\mathcal{S}}$ for the next iteration by finding the elements in $\boldsymbol{\mu}_{\alpha}$ with sufficiently large energy. Specifically, we set a small threshold $\varepsilon>0$ and compare each element of $\boldsymbol{\mu}$ with the threshold to obtain the estimated support as
@@
where the threshold $\varepsilon$ is chosen according to the noise power $1/\left\langle\kappa\right\rangle$. A good choice for $\varepsilon$ is 2 to 3 times the noise power $1/\left\langle\kappa\right\rangle$.
@@
Update Equation for $q\left(\kappa\right)$
The posterior distribution $q\left(\kappa\right)$ is given by
$q\left(\kappa\right)=\Gamma\left(\kappa;\check{c},\check{d}\right),$ (38)
@@
$\left\langle\rho_{k,q}\right\rangle=\dfrac{\check{a}_{k,q}}{\check{b}_{k,q}},\text{ }\left\langle s_{k,q}\right\rangle=\check{\lambda}_{k,q},\text{ }\left\langle\kappa\right\rangle=\dfrac{\check{c}}{\check{d}},$

### v2

@@ Abstract
ISAC, joint sensing and data recovery, blinear subspace variational Bayesian inference.
@@ I Introduction
Joint scattering environment sensing and channel estimation: Since multiple users in the same cell share a common scattering environment, the three-dimensional (3D) scatterers seen by different users have a partially overlapping structure. The work in [5, 6] revealed that MU channels exhibit a hidden joint sparsity in the angular-domain due to the shared scattering environment, which can be exploited to enhance the channel estimation performance. [3, 7] concentrated on the partially overlapping structure between radar targets and communication scatterers, and a joint target sensing and channel estimation scheme was designed based the joint sparsity of radar and communication channels. The above works only consider the oversimplified case when MU or ISAC channels share a partially overlapping support in the angular domain, while the shared location-domain information is not fully used. A recent work in [8] has proposed a two-dimensional (2D) location-domain sparse channel model to exploit the partially overlapping support of MU channels for the special case when all scatterers lies in a 2D area. However, the 3D location-domain sparse channel modeling is still not fully investigated. Moreover, the existing works perform joint scattering environment sensing and channel estimation based on pilot signals only. Nevertheless, data signals also provides much information about wireless channels that can help reduce pilot overhead.
@@
  • BiSVBI algorithm: We model the joint sensing and data recovery problem as a bilinear structured sparse signal recovery problem with a dynamic grid and imperfect factors. Then, we propose an expectation maximization based turbo BiSVBI (EM-Turbo-BiSVBI) algorithm to solve this problem efficiently, where the E-step performs Bayesian estimation of the location-domain sparse channels and the M-step refines the dynamic grid and learns the imperfect factors via gradient update. In particular, the E-step is based on a novel BiSVBI algorithm by minimizing the KL-divergence (KLD) between the variational posterior distribution and the true posterior associated with the bilinear model. To reduce the complexity, an independent distributed constraint is imposed on the variational posterior distribution of the sparse channels to avoid matrix inverse in the calculation of the posterior covariance matrix. Moreover, a subspace constrained matrix inverse followed by a gradient update is proposed to replace the high dimensional matrix inverse in the calculation of the posterior mean. Theoretical analysis and simulations show that BiSVBI greatly reduces the algorithm complexity with almost no sacrifice on the performance and convergence speed.
@@ II-B Uplink Signal Model for Joint Sensing and Data Recovery
where $\alpha_{k,0}$ and $\alpha_{k,l}$ are the complex gain of the LoS channel path and the $l$-th NLoS channel path, respectively, which are treated as random variables to facilitate Bayesian algorithm design. $f_{0}$ is the subcarrier spacing, $\tau_{o,k}$ denotes the timing synchronization error between user $k$ and the BS, $\tau_{0}\left(\boldsymbol{p}_{u,k}\right)=\frac{r_{l}}{c}$ is the propagation delay of the LoS path from user $k$ to the BS with $c$ denoting the speed of light,
@@ III-B Three-layer BGG Prior Model
where $\delta\left(\cdot\right)$ is the Dirac Delta function, $\Gamma\left(\rho;a,b\right)$ is a Gamma hyper-prior with shape parameter $a$ and rate parameter $b$. When $s_{k,q}=1$, $\alpha_{k,q}$ is a non-zero element and the corresponding variance $1/\rho_{k,q}$ is $\mathcal{O}\left(1\right)$. In this case, $a$ and $b$ should be chosen to satisfy $\tfrac{a}{b}=\mathbb{E}\left(\rho_{k,q}\right)=\mathcal{O}\left(1\right)$. When $s_{k,q}=0$, $\alpha_{k,q}$ is a zero element and the corresponding variance $1/\rho_{k,q}$ is close to zero. In this case, $\overline{a}$ and $\overline{b}$ should be chosen to satisfy $\tfrac{\overline{a}}{\overline{b}}=\mathbb{E}\left(\rho_{k,q}\right)\gg 1$. A typical value is $a=1$, $b=1$, $\overline{a}=1$, and $\overline{b}=10^{-5}$ [19].
@@ III-C Bilinear Sparse Bayesian Inference with Uncertain Parameters
where $\overline{\boldsymbol{v}}\triangleq\left\{\boldsymbol{\alpha},\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{s},\overline{\boldsymbol{s}},\gamma\right\}$ denotes the collection of all variables, and $p\left(\boldsymbol{\xi}\right)$ denotes the known prior distribution of $\boldsymbol{\xi}$. Once we obtain the MAP estimate of $\boldsymbol{\xi}^{\ast}$, we can obtain the minimum mean square error (MMSE) estimate of $\boldsymbol{\alpha}$ as $\boldsymbol{\alpha}^{*}=\int_{\boldsymbol{\alpha}}\boldsymbol{\alpha}p\left(\boldsymbol{\alpha}\mid\boldsymbol{y};\boldsymbol{\xi}^{\ast}\right)$, and the MAP estimate of $\boldsymbol{s}$ as $s_{k,q}^{\ast}=\arg\max_{s_{k,q}}p\left(s_{k,q}\mid\boldsymbol{y};\boldsymbol{\xi}^{\ast}\right),\forall k,q$. Moreover, the marginal posteriors $p\left(\boldsymbol{x}\mid\boldsymbol{y};\boldsymbol{\xi}\right)$ for the data can be used as soft information for digital demodulation and channel decoding.
@@ IV-B Turbo-BiSVBI-E Step
For each turbo iteration, Module A treats $\textrm{h}_{k,q}^{\text{A}}$ as the prior and performs the BiSVBI to calculate the approximate conditional posteriors $q\left(s_{k,q}\right)$’s. Then, Module A passes extrinsic messages to Module B by subtracting the prior information from posterior information in the log-probability domain, i.e.,
@@ IV-C1 Formulation of BiSVBI
The BiSVBI module generalizes the SC-VBI in [16] from a linear observation model to a bilinear model to calculate the approximate marginal posterior $q\left(\boldsymbol{v}\right)$ for fixed sensing parameters $\boldsymbol{\xi}$, where $\boldsymbol{v}\triangleq\left\{\boldsymbol{\alpha},\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{s},\gamma\right\}$. To be more specific, the turbo-iteration factor nodes $\textrm{h}_{k,q}^{\text{A}}\left(s_{k,q}\right),\forall k,q$ incorporate the prior information from Module B. Therefore, the following prior distribution is assumed in the BiSVBI module:
@@ IV-C2 Derivation of the BiSVBI Algorithm for $\mathcal{\mathscr{A}}_{\mathrm{VBI}}$
where $\mathbf{W}_{\alpha}=\textrm{Diag}\left(\left\langle\boldsymbol{\rho}\right\rangle\right)+\left\langle\gamma\right\rangle\left\langle\mathbf{A}_{\alpha}^{H}\mathbf{A}_{\alpha}\right\rangle$, $\mathrm{W}_{\alpha,k,q}$ is the $\left[\left(k-1\right)\left(Q+1\right)+q+1\right]$-th diagonal element of $\mathbf{W}_{\alpha}$, and $\boldsymbol{b}_{\alpha}=\left\langle\gamma\right\rangle\left\langle\mathbf{A}_{\alpha}^{H}\right\rangle\boldsymbol{y}$ with $\mathbf{A}_{\alpha}=\left[\mathbf{A}_{\alpha}^{p};\mathbf{A}_{\alpha}^{d}\right]$,
@@
where $\boldsymbol{\mu}_{x,t_{d},n}$ and $\boldsymbol{\sigma}_{x,t_{d},n}^{2}$ are the posterior mean and variance corresponding to $\boldsymbol{x}_{t_{d},n}$, $\mathbf{W}_{x,t_{d},n}\triangleq\textrm{Diag}\left(\boldsymbol{\sigma}_{t_{d},n}^{2}\right)+\left\langle\gamma\right\rangle\left\langle\mathbf{\Psi}_{n}^{H}\mathbf{\Psi}_{n}\right\rangle$ with $\boldsymbol{\sigma}_{t_{d},n}^{2}=\left[\sigma_{k,n}^{2}\right]_{k\in\mathcal{K}_{t_{d},n}^{d}}$, $\boldsymbol{b}_{x,t_{d},n}=\left\langle\gamma\right\rangle\left\langle\mathbf{\Psi}_{n}^{H}\right\rangle\boldsymbol{y}_{t_{d},n}^{d}$, and $\left[\mathrm{W}_{x,t_{d},k,n}\right]_{k\in\mathcal{K}_{t_{d},n}^{d}}=\textrm{Diag}\left(\mathbf{W}_{x,t_{d},n}\right)$.
Please refer to Appendix A for the detailed proof. Based on (25) and Lemma 2, the update equations of all variables are given below.
@@
$\boldsymbol{\mu}_{\alpha,\hat{\mathcal{S}}}^{s}=\left\langle\gamma\right\rangle\mathbf{W}_{\alpha,\hat{\mathcal{S}}}^{-1}\left\langle\mathbf{A}_{\alpha,\hat{\mathcal{S}}}^{H}\right\rangle\boldsymbol{y},$ (29)
@@
To improve the robustness of the algorithm against the support estimation error, we choose the best initial point $\boldsymbol{\mu}_{\alpha}^{0}$ from the following two choices: 1) $\boldsymbol{\mu}_{\alpha}^{s}$ obtained by the subspace constrained matrix inverse as in (29); 2) the posterior mean $\boldsymbol{\mu}_{\alpha}^{p}$ from the previous iteration, depending on which choice gives a lower value of the objective function $\varphi_{\alpha}\left(\boldsymbol{\mu}_{\alpha}^{0}\right)$. In the first few iterations, $\boldsymbol{\mu}_{\alpha}^{s}$ is expected to be a good choice because the subspace constrained matrix inverse can be used to accelerate the initial convergence speed, compared to the MM-based methods in [22, 8, 23] which completely avoid any matrix inverse. As the iterations go on, the posterior mean $\boldsymbol{\mu}_{\alpha}^{p}$ from the previous iteration becomes more and more accurate and it may also become a good initial point.
@@
$\boldsymbol{\mu}_{\alpha}^{\left(i\right)}$ $=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}-\epsilon_{\alpha}^{\left(i\right)}\nabla_{\boldsymbol{u}}\varphi_{\alpha}\left(\boldsymbol{u}_{\alpha}\right)\mid_{\boldsymbol{u}_{\alpha}=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}},$ (30)
@@
$\nabla_{\boldsymbol{u}}\varphi_{\alpha}\left(\boldsymbol{u}_{\alpha}\right)\mid_{\boldsymbol{u}_{\alpha}=\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}}=\mathbf{W}_{\alpha}\boldsymbol{\mu}_{\alpha}^{\left(i-1\right)}-\boldsymbol{b}_{\alpha},$ (31)
@@
Once we obtain the estimated posterior mean $\boldsymbol{\mu}_{\alpha}$, we can obtain an estimated support $\hat{\mathcal{S}}$ for the next iteration by finding the elements in $\boldsymbol{\mu}_{\alpha}$ with sufficiently large energy. Specifically, we set a small threshold $\varepsilon>0$ and compare each element of $\boldsymbol{\mu}_{\alpha}$ with the threshold to obtain the estimated support as
@@
where the threshold $\varepsilon$ is chosen according to the noise power $1/\left\langle\gamma\right\rangle$. A good choice for $\varepsilon$ is 2 to 3 times the noise power $1/\left\langle\gamma\right\rangle$.
@@
Update Equation for $q\left(\gamma\right)$
The posterior distribution $q\left(\gamma\right)$ is given by
$q\left(\gamma\right)=\Gamma\left(\gamma;\check{c},\check{d}\right),$ (38)
@@
$\left\langle\rho_{k,q}\right\rangle=\dfrac{\check{a}_{k,q}}{\check{b}_{k,q}},\text{ }\left\langle s_{k,q}\right\rangle=\check{\lambda}_{k,q},\text{ }\left\langle\gamma\right\rangle=\dfrac{\check{c}}{\check{d}},$